OpenStudy (anonymous):

8^-2= ??? A. 16 B. 64 C. -1/64 D. 1/64

OpenStudy (anonymous):

1/8^2=1/64

OpenStudy (anonymous):

\[8^{-2} = \frac{ 1 }{ 8^{2} }\]

OpenStudy (studygurl14):



OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

option D

OpenStudy (mathstudent55):

This is the definition of a negative exponent: \(a^{-n} = \dfrac{1}{a^n} \) Now just follow the definition and plug in your values.